OpenStudy (anonymous):

(2 - 3i)^2(2 + 3i)^2

OpenStudy (saifoo.khan):

169

OpenStudy (anonymous):

well that can be rewritten as ((2-3i)(2+3i))^2..now we know that (2-3i)(2+3i)=4-9i^2...difference between two squares rule in reverse. now we get (4-9i^2)^2..which is 16 - 72i^2 + 81i^4...16+72+81=169

OpenStudy (anonymous):

Yeap 169